OpenStudy (anonymous):

Solve 7W= -21. A. 147 B. -147 C. 3 D. -3

hero (hero):

7w = -21 w = -21/7 w = -3

OpenStudy (anonymous):

she seems to have problems solving for variables

hero (hero):

Well, it is a math problem, lol

OpenStudy (anonymous):

lol yeap, but she keeps asking fr the same type o f answers

OpenStudy (anonymous):

she obviously not understanding how to do it

hero (hero):

no need to state the obvious kid

OpenStudy (anonymous):

just saying obiously ....maybe explain better

hero (hero):

If you want, you can explain it in more detail. That's why they call it Open Study. More than one person can help.

OpenStudy (anonymous):

and she wont ask anymore

OpenStudy (anonymous):

yeap.

OpenStudy (anonymous):

If x =3 and y = -2 , then y4 - 4x = ? A. -28 B. -20 C. -4 D. 4

OpenStudy (anonymous):

(substitute)

OpenStudy (anonymous):

B..

OpenStudy (anonymous):

just plug in 3 for x and -2 for y and solve

OpenStudy (anonymous):

-2(4)-4(3)=